OpenStudy (anonymous):

simplify x+3/4x^2- 2x-4/9+7x/24

OpenStudy (espex):

\(x+\frac{3}{4}x^2-2x-\frac{4}{9}+\frac{7x}{24}\)?

OpenStudy (anonymous):

thanks :)

OpenStudy (espex):

I was just clarifying that that was the question you were asking. If that helped you solve it, then you are quite welcome. :D

OpenStudy (anonymous):

Sorry, I saw it wrong. Um, that isnt how it looks in the problem. its more like x-3 over 4x^2 minus 2x-4 over 9 plus 7x over 24. Does that make sense??

OpenStudy (whpalmer4):

@smc837 then you should either use the equation editor to write it that way, or use parentheses to correctly show your intended meaning. what @eSpeX wrote is what you have written, accounting for the operator precedence. what you should have written is (x+3)/(4x^2) - (2x-4)/9 + 7x/24 which is the equivalent to \[\frac{x+3}{4x^2} - \frac{2x-4}{9} + \frac{7x}{24}\]

OpenStudy (whpalmer4):

remember PEMDAS — you do multiplication and division before addition and subtraction, reading from left to right, so with what you wrote, we do the 3/4x^2 and the 4/9 and the 7x/24 before we do any addition.

OpenStudy (whpalmer4):

assume that if someone can interpret it differently, they will, and put in the necessary parentheses to force the correct understanding. unfortunately, going from a two-dimensional representation on the screen to a single line loses many of the visual cues that would make it clearer, so you have to put them back in with parentheses

OpenStudy (anonymous):

thank you for the explanation. It has really helped me. :)